OpenStudy (anonymous):

y = 3/(8+x). Calculate the 2nd and 3rd derivative. (show steps)

OpenStudy (anonymous):

First you have to calculate the first derivative so to do that use the Quotient Rule \[\frac{(8 + x)(0) - (3)(1)}{(8 + x)^2}\]

OpenStudy (anonymous):

I was told to use something called the Chain Rule.

OpenStudy (anonymous):

The chain rule is used only when you have a composite function like f(g(x)).. What you have here is a quotient \[\frac{f(x)}{g(x)}\]

OpenStudy (anonymous):

change to y = 3(8+x)^-1 and the chain rule becomes easier to see

OpenStudy (anonymous):

Yeah you can do it that way but its not really necessary

OpenStudy (anonymous):

I did the quotient rule, then I did it again, and the answer was wrong.

OpenStudy (anonymous):

so y' = (-1)(3)(8+x)^-1-1 =-3(8+x)^-2

OpenStudy (anonymous):

and use the same method for y'' and y'''

OpenStudy (anonymous):

The first derivative is \[- \frac{3}{(8 + x)^2}\]

OpenStudy (anonymous):

How did you get that as your derivative?

OpenStudy (anonymous):

now it would be easier to use the chain rule so rewrite it as \[-3(8 + x)^{-2}\]

OpenStudy (anonymous):

so the second derivative will be \[6(8 + x)^{-3} * 1\] or \[\frac{6}{(8 + x)^{3}}\]

OpenStudy (anonymous):

and the third derivative would be \[- \frac{18}{(8+x)^{4}}\]